# HawkDrive: A Transformer-driven Visual Perception System for Autonomous Driving in Night Scene

### v1

@@ I-A Motivation
Visual perception of self-driving vehicles in low light conditions remain a challenge. Insufficient and patchy illumination in the night driving environment can cause considerable interference to cameras, thus adversely affecting the performance of neural network[1].
Besides, data-driven monocular depth estimation for self-driving are susceptible to adversarial conditions such as image noise, occlusions, variations in lighting, and surface reflections, while the stereo camera is demonstrated with more sufficient accuracy comparing to the monocular one [2], [3]. Based on the data processing from stereo camera, a flexible edge computing method should be developed since it performs capturing and processing data as close to its source or end user as possible to empower the perception system with low latency for autonomous driving. For more reliable downstream task performance, as the learning-based perception models are developed less vigorously, it is also necessary to deploy state-of-the-art neural networks on the edge computation devices for inference and evaluation.
[…]
@@ I-B Problem Statement
Errors in depth estimation and image noise caused by unfavorable sensing conditions are critical factors of failure in perception, navigation, and planning.[4]. Passive sensors, such as RGB cameras, can acquire noisy, low-dynamic range and blurry images. Therefore, a versatile pipeline for post-processing raw sensor data is required. Moreover, satisfactory processing speed is needed in order to adapt to various driving situations. In particular, it is challenging for cameras to discover full details in low-light conditions. A few prior works have been proposed to enhance night-time camera data, such as TodayGAN [5], EnlightenGAN [6]. Nevertheless, these techniques are not able to manage high dynamic range scenes, as they are unable to maintain enough details for driving scenes that have a variety of dynamic ranges of sight and illumination conditions due to the non-uniform boosting areas of the entire frame [7].
@@ I-D Contribution
The image data obtained from the stereo cameras require accurate hardware synchronization and low latency. Thus, hardware trigger cables and Nvidia Jetson Xavier AGX are partnered within an adjustable in-car structure for experimental validation combining 3D-printed computation device base, camera holders, and pneumatic support in Fig. 2.
@@
Figure 2: Decomposition of the HawkDrive structure.
Figure 3: HawkDrive installed in the self-driving car.
@@ II SYSTEM OVERVIEW
Figure 4: System overview. Raw images are captured and published towards image processing node, where the rectification and other necessary processing are performed. Resize node receives the preprocessed images and passes them to downstream modules. Here is the workflow for night driving scene, ROS2 messages of resized image are published to low light enhancement and segmentation modules to perform semantic knowledge based enhancement. Then the enhanced images are conveyed to other modules for corresponding tasks.
@@ II-A Hardware
Two global-shutter The Imaging Source DFK 33UX250 cameras with 2.3 mm machine vision lens Computar T2314 FICS, CMOS Pregius Sony IMX250 sensor are used as capture modules which can provide ($2,448\times 2,048$) 5MP resolution and 12 bit dynamic range. Two global shutter cameras with a 67.12 cm baseline capture images of the entire sensor at once, ensuring that all pixels are exposed simultaneously [20]. Utilizing a synchronized pair of global shutter cameras eliminates the rolling shutter effect and ensures accurate timing of events.
@@
Therefore, the combination of the Nvidia Jetson Xavier AGX and stereo vision (see Fig. 3) is able to offer a promising solution for autonomous driving scenes [21].
@@ II-C Pipeline
Initiating the camera ROS2 driver, the stereo images from the left and right cameras are saved as dataset for stereo calibration preparation. Then, using stereo calibration algorithm, the stereo calibration results are obtained. For processing the calibration matrix, two methods are developed. First, the calibration matrix is integrated into ROS2 node configuration which will be done by filling the Camera_Info messages. Second, the calibration results are integrated into ROS2 package with OpenCV [23].
Receiving the current from Jetson pins through an amplifier, the two cameras are connected via trigger cables. The pulse output is controlled by the trigger module to control the camera frame rate and exposure time input. Then a ROS2 package for receiving hardware synchronized images from left and right camera is built, transforming the image data to ROS2 messages and publishing the synchronized messages via ROS2 nodes. For downstream processing, the message subscribers and synchronizers are respectively established. Through the built driver, rectified images from stereo camera are output, while they are then subscribed by the low light enhancement module, depth prediction and semantic segmentation module. All the modules work in customized docker containers [24], [25] and the whole system overview can be concluded as Fig. 4.
@@ III-A Signal-to-Noise-Ratio-aware Low Light Enhancement
Image noise is treated as discontinuous transition between adjacent pixels in the spatial domain. The SNR map obtained from the distance between the night image and an associated gray-scale image is used to guide the fusion of long-range (transformer processing results) and short-range (convolutional neural network processing results) features (i.e. the fusion of local information and global information).
Our resize node resizes the raw images to a resolution of $600\times 400$, which are then sent to the SNR-aware enhancement module to produce the enhanced images. In Fig. 5, our dataset’s results show a considerable improvement in the night scene, while the details of vehicles, pedestrians, and roads were preserved.
@@ III-B Global Matching Depth Estimation
For higher real-time performance, the Unimatch module subscribes to the resized $600\times 400$ image messages. Setting the module output as stereo matching, the results from our dataset is shown in Fig. 6. This figure shows the filtered depth at each pixel, by averaging the prediction over a small patch centered at this pixel.
Ranftl et al. [19] proposed a dense prediction transformer (DPT) specifically for dense prediction tasks, while the monocular depth estimation can be treated as a dense regression problem. It showed a performance improvement compared to the fully convolutional neural network, so the monocular depth estimation of DPT module is adapted for the cases when one of the stereo cameras has malfunction. The predicted depth of DPT module output is shown in Fig. 7.
(a) Input camera image.
(b) Output depth image.
Figure 7: DPT depth estimation module processing.
@@ III-C SegFormer
[…]
@@
(a) Input camera image.
(b) Output segmented image.
Figure 9: SegFormer module processing.
@@ IV-A Experimental Setup
Figure 10: Experimental Self-Driving Car.
With onboard servers and power supply of the experimental car, shown in Fig. 10, the test dataset was acquired along the same driving routine from daytime and nighttime driving scenes, while two cameras were hardware triggered and Jetson Xavier AGX was switched to power mode with camera ROS2 driver and each module running respectively. The sequential image samples of the recorded dataset are shown in Fig. 11.
For synchronization monitoring, the ROS2 message timestamps could be checked along with the ROS2 logger information and callbacks. For module running checking, callback functions with inference time and frames per second (FPS) could demonstrate the system performance. To adapt to the day and night light conditions, the camera parameters such as exposure time, gain, frame rate, etc. were tuned from the configuration files in order to capture the necessary scene information.
Figure 11: Captured sequential images of the recorded dataset: night driving scene (above), day driving scene (below).
@@ IV-B Enhancement Results Comparison
First, the performance was evaluated on our recorded dataset. Scale Invariant Feature Transform (SIFT) presented by D. Lowe [29] has significant applications on object recognition and robot navigation. Feature analysis of 50 images from different critical driving scenes containing pedestrians, vehicles and traffic signs, etc. iconically was performed based on SIFT algorithms. The results in Fig. 13 showed that on average, 148.1 more key points of objects in the images can be detected after enhancement.
For depth estimation performance, visual comparison between the night scene and the enhanced scene is provided in Fig. 14, which demonstrates that due to the poor illumination, the detected targets did not display necessary contour for recognition after the processing. However, the processed enhanced image improved the contour intactness. From our experimental self-driving car, Velodyne Ultra Puck (VLP-32) LiDAR is used to provide the ground truth of distance measurements. Obtaining the ground truth depth from annotated LiDAR point clouds shown in Fig. 12, the errors between LiDAR and stereo camera depth estimation were reduced $27.16\%$ after the enhancement.
@@
[…]
@@
Figure 15: Waypoints’ translation errors of trajectory estimation on KITTI Odometry sequence 00 under day, night and enhanced conditions. The mean errors of night scene was $107.176$ m, while the mean errors of enhanced scene from our enhancement method reached $44.053$ m comparing to $46.470$ m in day condition.
TABLE I: Average pixel accuracy comparison of our recorded dataset samples. Among 50 samples of critical scenes, here is the showcase of random 6 samples out of 50 samples.

### v2

@@ I-A Motivation
Visual perception of self-driving vehicles in low light conditions remains a challenge. Insufficient and patchy illumination in the night driving environment can cause considerable interference to cameras, thus adversely affecting the performance of neural network [1].
Besides, data-driven monocular depth estimation for self-driving is susceptible to adversarial conditions such as image noise, occlusions, variations in lighting, and surface reflections, while the stereo camera is demonstrated with more sufficient accuracy compared to the monocular one [2], [3]. Based on the data processing from stereo camera, a flexible edge computing method should be developed since it performs capturing and processing data as close to its source or end user as possible to empower the perception system with low latency for autonomous driving. For more reliable downstream task performance, as the learning-based perception models are developed less vigorously, it is also necessary to deploy state-of-the-art neural networks on the edge computation devices for inference and evaluation.
(a) Semantic segmentation of night image.
(b) Semantic segmentation of enhanced image.
@@ I-B Problem Statement
Errors in depth estimation and image noise caused by unfavorable sensing conditions are critical factors of failure in perception, navigation, and planning [4]. Passive sensors, such as RGB cameras, can acquire noisy, low-dynamic range and blurry images. Therefore, a versatile pipeline for post-processing raw sensor data is required. Moreover, a satisfactory processing speed is needed in order to adapt to various driving situations. In particular, it is challenging for cameras to discover full details in low-light conditions. A few prior works have been proposed to enhance night-time camera data, such as TodayGAN [5], EnlightenGAN [6]. Nevertheless, these techniques are not able to manage high dynamic range scenes, as they are unable to maintain enough details for driving scenes that have a variety of dynamic ranges of sight and illumination conditions due to the non-uniform boosting areas of the entire frame [7].
@@ I-D Contribution
The image data obtained from the stereo cameras requires accurate hardware synchronization and low latency. Thus, hardware trigger cables and Nvidia Jetson Xavier AGX are partnered within an adjustable in-car structure for experimental validation combining 3D-printed computation device base, camera holders and pneumatic support.
@@ II System Overview
Figure 2: System overview. Here is the workflow for night driving scene, ROS2 messages of resized image are published to low light enhancement and segmentation modules to perform semantic knowledge based enhancement. Then the enhanced images are conveyed to other modules for corresponding tasks.
@@ II-A Hardware
Two global-shutter The Imaging Source DFK 33UX250 cameras with 2.3 mm machine vision lens Computar T2314 FICS, CMOS Pregius Sony IMX250 sensor are used as capture modules which can provide a ($2,448\times 2,048$) 5MP resolution and 12 bit dynamic range. Two global shutter cameras with a 67.12 cm baseline capture images of the entire sensor at once, ensuring that all pixels are exposed simultaneously [20]. Utilizing a synchronized pair of global shutter cameras eliminates the rolling shutter effect and ensures accurate timing of events.
@@
Therefore, the combination of the Nvidia Jetson Xavier AGX and stereo vision is able to offer a promising solution for autonomous driving scenes [21].
@@ II-C Pipeline
Initiating the camera ROS2 driver, the stereo images from the left and right cameras are saved as dataset for stereo calibration preparation. Then, using a stereo calibration algorithm, the stereo calibration results are obtained. For processing the calibration matrix, two methods are developed. First, the calibration matrix is integrated into the ROS2 node configuration which will be done by filling the Camera_Info messages. Second, the calibration results are integrated into the ROS2 package with OpenCV [23].
Receiving the current from Jetson pins through an amplifier, the two cameras are connected via trigger cables. The pulse output is controlled by the trigger module to control the camera frame rate and exposure time input. Then a ROS2 package for receiving hardware synchronized images from left and right camera is built, transforming the image data to ROS2 messages and publishing the synchronized messages via ROS2 nodes. For downstream processing, the message subscribers and synchronizers are respectively established. Through the built driver, rectified images from the stereo camera are output, while they are then subscribed by the low light enhancement module, depth prediction and semantic segmentation module. All the modules work in customized docker containers [24], [25] and the whole system overview can be concluded as Fig. 2.
@@ III-A Signal-to-Noise-Ratio-aware Low Light Enhancement
Image noise is treated as a discontinuous transition between adjacent pixels in the spatial domain. The SNR map obtained from the distance between the night image and an associated gray-scale image is used to guide the fusion of long-range (transformer processing results) and short-range (convolutional neural network processing results) features (i.e. the fusion of local information and global information).
Our resize node resizes the raw image to a resolution of $600\times 400$, which is then sent to the SNR-aware enhancement module to produce the enhanced image. In Fig. 4, our dataset’s results show a considerable improvement in the night scene, while the details of vehicles, pedestrians, and roads were preserved.
@@ III-B Global Matching Depth Estimation
For higher real-time performance, the Unimatch module subscribes to the resized $600\times 400$ image messages. Setting the module output as stereo matching, the results from our dataset are shown in Fig. 5. This figure shows the filtered depth at each pixel, by averaging the prediction over a small patch centered at this pixel.
Ranftl et al. [19] proposed a dense prediction transformer (DPT) specifically for dense prediction tasks, while the monocular depth estimation can be treated as a dense regression problem. It showed a performance improvement compared to the fully convolutional neural network, so the monocular depth estimation of DPT module is adapted for the cases when one of the stereo cameras has malfunction.
@@ III-C SegFormer
Meanwhile, the SegFormer module subscribes to raw image input and publishes segmentation map to low light enhancement module as prior knowledge input to refine the boosting results. In Fig. 7, night image $I_{n}$ is conveyed to segmentation network and enhancement network to obtain segmentation map $I_{seg}$ and fusion feature $F_{f}$. The attention map $A$ is computed as follows,
@@ IV-A Experimental Setup
With onboard servers and power supply of the experimental car, shown in Fig. 3, the test dataset was acquired along the same driving routine from daytime and nighttime driving scenes, while two cameras were hardware triggered and Jetson Xavier AGX was switched to power mode with camera ROS2 driver and each module running respectively.
For synchronization monitoring, the ROS2 message timestamps can be checked along with the ROS2 logger information and callbacks. For module running checking, callback functions with inference time and frames per second (FPS) could demonstrate the system performance. To adapt to the day and night light conditions, the camera parameters such as exposure time, gain, frame rate, etc. were tuned from the configuration files in order to capture the necessary scene information.
@@ IV-B Enhancement Results Comparison
First, the performance was evaluated on our recorded dataset. Scale Invariant Feature Transform (SIFT) presented by D. Lowe [29] has significant applications in object recognition and robot navigation. Feature analysis of 50 images from different critical driving scenes containing pedestrians, vehicles and traffic signs, etc. iconically was performed based on SIFT algorithms. The results in Fig. 8 showed that on average, 148.1 more key points of objects in the images can be detected after enhancement.
For depth estimation performance, visual comparison between the night scene and the enhanced scene is provided in Fig. 9, which demonstrates that due to the poor illumination, the detected targets did not display the necessary contour for recognition after the processing. However, the processed enhanced image improved the contour intactness. On our experimental self-driving car, Velodyne Ultra Puck (VLP-32) LiDAR is used to provide the ground truth of distance measurements. Obtaining the ground truth depth from annotated LiDAR point clouds shown in Fig. 6, the errors between LiDAR and stereo camera depth estimation were reduced $27.16\%$ after the enhancement.
@@
(b) Output depth of the night image.
(c) Output depth of the enhanced image.
@@
Figure 10: Waypoints’ translation errors of trajectory estimation on KITTI Odometry sequence 00 under day, night and enhanced conditions. The mean errors of night scene was $107.176$ m, while the mean errors of enhanced scene from our enhancement method reached $44.053$ m compared to $46.470$ m in day condition.
TABLE I: Average pixel accuracy comparison of our recorded dataset samples. Among 50 samples of critical scenes with intensive noise input, here is the showcase of 6 random samples out of 50 samples.